OpenStudy (anonymous):

solve the equation by graphing 5x-9y=-54 9y-5x=54

OpenStudy (amistre64):

umm..usually by graphing it :) and finding the points of intersection. It is the most unreliable way to get a good answer tho, but it gives you an idea on what a solution means.

OpenStudy (amistre64):

I can see that your equations are perpendicular lines that cross....

OpenStudy (amistre64):

nonono....they are parallel lines.... always sem to get the mixed ;)

OpenStudy (amistre64):

5x -9y = -54 -5x +9y = 54 <- if we multiply this by -1 we get: -1(5x -9y = -54) we are back to your same equation

OpenStudy (amistre64):

the answer is either all points are solutions or no points :)

OpenStudy (amistre64):

all points is my vote...

OpenStudy (anonymous):

I agree. There are integer solutions: x=9n, y=5n+6, and \[n \in \mathbb{Z}\]

OpenStudy (anonymous):

before yo graphng it you have to change it to y inetersept form then make a atable of values then you get your corrdintaes together then you graph

OpenStudy (anonymous):

right

OpenStudy (anonymous):

right

OpenStudy (anonymous):

not sure y it keeps typing that in

OpenStudy (anonymous):

\[5x-9y=-54\]

OpenStudy (anonymous):

what

OpenStudy (anonymous):

are you a student

OpenStudy (anonymous):

you there

OpenStudy (anonymous):

yes i am sry

OpenStudy (anonymous):

the problem i am working on now is y= -x -8 9x-8y=-21

OpenStudy (anonymous):

are you good at writing essays

OpenStudy (anonymous):

so so what school do u go to

OpenStudy (anonymous):

ms 137

OpenStudy (anonymous):

what school

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

u don't know what school u go to?

OpenStudy (anonymous):

u don't know what school u go to?